if [x] is the integral part of x, then solve


[4-[x]^2}-[x]^2=2

Grade:12

1 Answers

Swapnil Saxena
12 years ago

If [x] is the greatest integer function for x, the x can belong to [1,2) and [-1,0}

Let [x] = I where I is an integer , then (4 - I^2)-I^2=2

Then 4-2I^2=2

=> I^2=-1

=> I = -1,1

if I = [x] then posiible solution are [1,2)and [-1,0).